# FAB-PPI: Frequentist, Assisted by Bayes, Prediction-Powered Inference

### v1

@@ 1 Introduction
Statistical inference crucially relies on the availability of high-quality labelled data to draw actionable conclusions. As the scale of machine learning models keeps growing, their increasingly accurate predictions become a tempting alternative to labelled data in fields where the latter are traditionally scarce, such as proteomics (Jumper et al., 2021). However, blindly using potentially biased predictions as a surrogate for labelled data voids the statistical validity of the conclusions drawn. To address this, prediction-powered inference (Angelopoulos et al., 2023b) provides a general framework for statistical inference in the presence of a large number of black-box predictions by combining them with a smaller number of labelled observations, which are used to correct for the discrepancy between the predictions and the true labels. The estimators and confidence intervals (CIs) resulting from PPI are statistically valid regardless of the machine learning model used. Moreover, when the predictions are good, PPI results in more accurate estimates and shorter CIs than traditional methods that rely solely on labelled data.
@@
For instance, if $\mathcal{L}_{\theta}(x,y)=(\theta-y)^{2}/2$ is the squared loss, then $\theta^{\star}=\mathbb{E}[Y]$. We assume that we have $n$ labelled observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$ iid from $\mathbb{P}$ and $N$ unlabelled observations $\{\widetilde{X}_{i}\}_{i=1}^{N}$ iid from $\mathbb{P}_{X}$, which are also independent of the labelled data. The number of unlabelled observations is typically much larger than the number of labelled ones, $N\gg n$. Additionally, we are provided with a machine learning prediction rule $f$, that can be used to predict an output $f(x)$ at any input $x$. PPI aims to obtain an estimator $\widehat{\theta}$ and a $1-\alpha$ CI $\mathcal{C}_{\alpha}^{\mathrm{pp}}$ for $\theta^{\star}$, which take advantage of $f$. Under mild assumption, $\theta^{\star}$ can be expressed as the solution to
@@
In this setting, $m_{\theta}$ represents a measure of fit of the predictor, whereas $\Delta_{\theta}$, called the rectifier, accounts for the discrepancy between the predicted outputs $f(X)$ and the true outputs $Y$. In the case of the squared loss, $\Delta_{\theta}=\mathbb{E}[f(X)-Y]$ does not depend on $\theta$, but this is not true in general. By estimating the two quantities $m_{\theta}$ and $\Delta_{\theta}$, Angelopoulos et al. (2023a) derive an estimator and a CI for $\theta^{\star}$, which use both labelled and unlabelled data. The resulting CI is shorter than the classical confidence interval based solely on the labelled data when $N\gg n$ and $f$ is accurate because, in this case, $m_{\theta}$ can be estimated with low variance using the unlabelled data, while $\Delta_{\theta}$ is close to zero.
Standard PPI employs off-the-shelf estimation and CI procedures for $\Delta_{\theta}$, which do not take advantage of any prior knowledge on the quality of the machine learning model $f$. However, in many applications, we expect the latter’s predictions to be (i) usually very good, but (ii) sometimes prone to large errors and hallucinations. We propose to encode such an inductive bias with a horseshoe prior $\pi_{\theta}$ on $\Delta_{\theta}$ (Carvalho et al., 2010), which accommodates the aforementioned properties by exhibiting (i) an infinitely tall spike at the origin, and (ii) Cauchy-like tails at infinity. In order to construct valid confidence regions for $\Delta_{\theta}$ using the horseshoe prior, we resort to the frequentist-assisted by Bayes (FAB) framework (Pratt, 1961, 1963; Yu & Hoff, 2018). This approach provides confidence regions such that their expected length is lower for rectifiers $\Delta_{\theta}$ that have high probability under $\pi_{\theta}$, and larger otherwise. While the resulting confidence regions have exact coverage for any prior $\pi_{\theta}$, the horseshoe prior is particularly well-suited for PPI. Being concentrated around the origin, it produces shorter confidence regions when the predictions are good, i.e. $||\Delta_{\theta}||\simeq 0$. At the same time, its heavy tails ensure robustness when the predictions are poor. Indeed, as shown by Cortinovis & Caron (2024), if $||\Delta_{\theta}||\gg 0$, the FAB procedure with the horseshoe prior reverts to the traditional CI based on the sample mean.
@@
  • Modular, as it can be used in conjunction with power tuning (Angelopoulos et al., 2023a).
@@ 2 Related work
PPI (Angelopoulos et al., 2023b) was introduced to obtain shorter CIs for the parameters of interest by leveraging machine learning predictions in semi-supervised settings. PPI has since been extended in multiple directions. PPI++ (Angelopoulos et al., 2023a) proposes a different, loss-based formulation of PPI, leading to a more computationally efficient procedure, along with an additional power-tuning parameter to enhance PPI’s performance. Stratified PPI (Fisch et al., 2024) improves upon PPI by employing a data stratification strategy. Cross PPI (Zrnic & Candès, 2024b) demonstrates how the training of $f$ can be included in the PPI pipeline. Active statistical inference (Zrnic & Candès, 2024a) applies an active learning approach to select which inputs from the unlabelled set should be labelled. Closer to our work, Bayesian PPI (Hofer et al., 2024) considers an alternative PPI estimator motivated by Bayesian ideas. However, their approach provides Bayesian credible intervals, which do not offer frequentist guarantees. Additionally, their approach achieves similar experimental performance to PPI, while we demonstrate that FAB-PPI may significantly improve upon PPI.
As discussed in Angelopoulos et al. (2023b, a), PPI has close ties with control variates for variance reduction (Glasserman, 2003, §4.1). In the case of mean estimation, the form of the PPI estimator is similar to the one proposed by Zhang et al. (2019). PPI is also related to work in semiparametric inference with missing data (Robins & Rotnitzky, 1995).
@@ 3.1 Control variates
The method of control variates is a standard variance reduction technique in Monte Carlo approximation (Glasserman, 2003, §4.1). For simplicity, we present the method in the scalar case, but extensions to the multivariate setting are available. Let $(Z,Y)$ be a pair of real-valued random variables, and assume we are interested in estimating $\mathbb{E}[Y]$ based on an iid sample $\{(Z_{i},Y_{i})\}_{i=1}^{n}$. Assuming $\mu=\mathbb{E}[Z]$ is known, one defines the control variate estimator (CVE)
@@
where $\lambda\in\mathbb{R}$ is a tuning coefficient and $\overline{Z}$ and $\overline{Y}$ are the sample means of $(Z_{i})$ and $(Y_{i})$, respectively. The centered random variable $Z_{i}-\mu$ serves as a control variate to estimate $\mathbb{E}[Y]$. The CVE is a consistent and unbiased estimator of $\mathbb{E}[Y]$ with $\mathrm{var}(\widehat{Y}^{\mathrm{cv}}_{\lambda})=(\mathrm{var}(Y)-2\lambda\mathrm{cov}(Z,Y)+\lambda^{2}\mathrm{var}(Z))/n$, while $\mathrm{var}(\overline{Y})=\mathrm{var}(Y)/n$. Therefore, CVE achieves smaller variance whenever $\lambda<\frac{2\mathrm{cov}(Z,Y)}{\mathrm{var}(Z)}$, and the optimal coefficient is given by $\lambda^{*}=\mathrm{cov}(Z,Y)/\mathrm{var}(Z)$. In this case, $\mathrm{var}(\widehat{Y}^{\mathrm{cv}}_{\lambda^{*}})=(1-\rho_{Z,Y}^{2})\mathrm{var}(\overline{Y})$, where $\rho_{Z,Y}$ is the correlation between $Z$ and $Y$. The more correlated $Z$ and $Y$, the larger the variance reduction. By plugging the estimator
@@
as $n\to\infty$, where $s$ is the sample standard deviation of $\{(Y_{i}-\widehat{\lambda}Z_{i})\}_{i=1,\ldots,n}$. Hence, $\widehat{Y}^{\mathrm{cv}}_{\widehat{\lambda}}\pm z_{1-\alpha/2}s/\sqrt{n}$ is an asymptotically valid $1-\alpha$ CI for $\mathbb{E}[Y]$, whose asymptotic width is $2z_{1-\alpha/2}\sqrt{1-\rho_{ZY}^{2}}\sqrt{\mathrm{var}(Y)}/\sqrt{n}$.
@@ 3.2 Prediction-powered inference
PPI (Angelopoulos et al., 2023b) defines an estimator $\widehat{\theta}$ and a CI $\mathcal{C}_{\alpha}^{\mathrm{pp}}$ for a parameter of interest $\theta^{\star}$ satisfying Equation 2. In particular, let $\widehat{m}_{\theta}$ and $\widehat{\Delta}_{\theta}$ be some estimators of $m_{\theta}$ and $\Delta_{\theta}$. Using Equation 2, the estimator $\widehat{\theta}$ is defined as the solution, in $\theta$, to the equation
@@
where $+$ denotes the Minkowski sum. Typical choices for $\widehat{m}_{\theta}$ and $\mathcal{T}_{\alpha-\delta}$ are the sample mean of the unlabelled data,
@@
Angelopoulos et al. (2023b) propose to use the sample mean
$\widehat{\Delta}^{\mathrm{\texttt{PP}}}_{\theta}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(\mathcal{L}_{\theta}^{\prime}(X_{i},Y_{i})-\mathcal{L}_{\theta}^{\prime}(X_{i},f(X_{i}))\right)$ (10)
@@
Angelopoulos et al. (2023a) extend standard PPI by introducing an additional control variate parameter $\lambda$, which they call power tuning parameter. The chosen $\widehat{m}_{\theta}$ is still the sample mean (9), while $\widehat{\Delta}^{\mathrm{\texttt{PP+}}}_{\theta}$ now takes the control variate form
@@
where $\widehat{\lambda}$ is estimated from the data. In this case, the centered control variate is $\mathcal{L}_{\theta}^{\prime}(X_{i},f(X_{i}))-\widehat{m}_{\theta}$, which depends only on the machine learning predictions. For the squared loss, we obtain
@@
While the definition of the PPI confidence interval (8) allows for merging any CIs $\mathcal{R}_{\delta}$ and $\mathcal{T}_{\alpha-\delta}$ for $\Delta_{\theta}$ and $m_{\theta}$, in practice these are often chosen to be CLT-based CIs that, once combined into $\mathcal{C}_{\alpha}$ give exact asymptotic coverage,
$\liminf_{n,N\to\infty}\mathrm{Pr}(\theta^{*}\in\mathcal{C}_{\alpha})\geq 1-\alpha.$
@@ 3.3 Bayes-optimal confidence regions
[…]
@@
[…]
@@ 4 FAB-PPI
As in PPI, we use the sample mean (9) as the estimator of $m_{\theta}$. For $\Delta_{\theta}$, we start by considering a consistent estimator $\widehat{\Delta}_{\theta}$, such as the sample mean (10) used in PPI, or the control variate estimator (12) used in PPI++. Throughout this section, we assume that 3.1 is satisfied. That is, a CLT holds for $\widehat{m}_{\theta}$ and $\widehat{\Delta}_{\theta}$ with respect to some estimators $(\widehat{\sigma}^{f}_{\theta})^{2}$ and $\widehat{\sigma}^{2}_{\theta}$ of $\mathrm{var}(\widehat{m}_{\theta})$ and $\mathrm{var}(\widehat{\Delta}_{\theta})$, respectively. In this setting, let $\pi_{0}(\Delta_{\theta};\tau_{n})$ be a prior on $\Delta_{\theta}$ with scale parameter $\tau_{n}$, which may depend on the labelled data through $\widehat{\sigma}_{\theta}$. Denote by $\ell(w;\sigma,\tau)$ the log-marginal likelihood, evaluated at $w$, of a Gaussian likelihood model with mean $\Delta$ and variance $\sigma^{2}$ under the prior $\pi_{0}(\Delta;\tau)$,
[…]

### v2

@@ 1 Introduction
Statistical inference crucially relies on the availability of high-quality labelled data to draw actionable conclusions. As the scale of machine learning models keeps growing, their increasingly accurate predictions become a tempting alternative to labelled data in fields where the latter are traditionally scarce, such as proteomics (Jumper et al., 2021). However, blindly using potentially biased predictions as a surrogate for labelled data voids the statistical validity of the conclusions drawn. To address this, prediction-powered inference (Angelopoulos et al., 2023a) provides a general framework for statistical inference in the presence of a large number of black-box predictions by combining them with a smaller number of labelled observations, which are used to correct for the discrepancy between the predictions and the true labels. The estimators and confidence intervals (CIs) resulting from PPI are statistically valid regardless of the machine learning model used. Moreover, when the predictions are good, PPI results in more accurate estimates and shorter CIs than traditional methods that rely solely on labelled data.
@@
For instance, if $\mathcal{L}_{\theta}(x,y)=(\theta-y)^{2}/2$ is the squared loss, then $\theta^{\star}=\mathbb{E}[Y]$. We assume that we have $n$ labelled observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$ iid from $\mathbb{P}$ and $N$ unlabelled observations $\{\widetilde{X}_{i}\}_{i=1}^{N}$ iid from $\mathbb{P}_{X}$, which are also independent of the labelled data. The number of unlabelled observations is typically much larger than the number of labelled ones, $N\gg n$. Additionally, we are provided with a machine learning prediction rule $f$, that can be used to predict an output $f(x)$ at any input $x$. PPI aims to obtain an estimator $\widehat{\theta}$ and a $(1-\alpha)$ confidence interval $\mathcal{C}_{\alpha}^{\mathrm{pp}}$ for $\theta^{\star}$, which take advantage of $f$. Under mild assumptions, $\theta^{\star}$ can be expressed as the solution to
@@
In this setting, $m_{\theta}$ represents a measure of fit of the predictor, whereas $\Delta_{\theta}$, called the rectifier, accounts for the discrepancy between the predicted outputs $f(X)$ and the true outputs $Y$, effectively quantifying prediction quality. For example, under the squared loss, $\Delta_{\theta}=\mathbb{E}[f(X)-Y]$ and a good predictor $f$ is one such that $\Delta_{\theta}$ is close to zero, i.e. $f(x)\simeq\mathbb{E}[Y|X=x]$. Note that, while in this case $\Delta_{\theta}$ does not depend on $\theta$, this is not true in general.
By estimating the two quantities $m_{\theta}$ and $\Delta_{\theta}$, Angelopoulos et al. (2023b) derive an estimator and a CI for $\theta^{\star}$, which use both labelled and unlabelled data. The resulting CI is shorter than the classical confidence interval based solely on the labelled data when $N\gg n$ and $f$ is accurate because, in this case, $m_{\theta}$ can be estimated with low variance using the unlabelled data, while $\Delta_{\theta}$ is close to zero.
Standard PPI employs off-the-shelf estimation and CI procedures for $\Delta_{\theta}$, which do not take advantage of any prior knowledge on the quality of the machine learning model $f$. However, in many applications, we expect the latter’s predictions to be (i) usually very good, but (ii) sometimes prone to large errors and hallucinations. We propose to encode such an inductive bias with a horseshoe prior $\pi_{\theta}$ on $\Delta_{\theta}$ (Carvalho et al., 2010), which accommodates the aforementioned properties by exhibiting (i) an infinitely tall spike at the origin, and (ii) Cauchy-like tails at infinity. In order to construct valid confidence regions for $\Delta_{\theta}$ using the horseshoe prior, we resort to the frequentist-assisted by Bayes (FAB) framework (Pratt, 1961, 1963; Yu & Hoff, 2018). This approach provides confidence regions such that their expected length is lower for rectifiers $\Delta_{\theta}$ that have high probability under $\pi_{\theta}$, and larger otherwise. While the resulting confidence regions have exact coverage for any prior $\pi_{\theta}$, the horseshoe prior is particularly well-suited for PPI. Being concentrated around the origin, it produces shorter confidence regions when the predictions are good, i.e. $\Delta_{\theta}||\simeq 0$. At the same time, its heavy tails ensure robustness when the predictions are poor. Indeed, as shown by Cortinovis & Caron (2024), if $\Delta_{\theta}||\gg 0$, the FAB procedure with the horseshoe prior reverts to the traditional CI based on the sample mean.
@@
  • Modular, as it can be used in conjunction with power tuning (Angelopoulos et al., 2023b).
@@ 2 Related Work
PPI (Angelopoulos et al., 2023a) was introduced to obtain shorter CIs for the parameters of interest by leveraging machine learning predictions in semi-supervised settings. PPI has since been extended in multiple directions. PPI++ (Angelopoulos et al., 2023b) proposes a different, loss-based formulation of PPI, leading to a more computationally efficient procedure, along with an additional power tuning parameter to enhance PPI’s performance. Stratified PPI (Fisch et al., 2024) improves upon PPI by employing a data stratification strategy. Cross PPI (Zrnic & Candès, 2024b) demonstrates how the training of $f$ can be included in the PPI pipeline. Active statistical inference (Zrnic & Candès, 2024a) applies an active learning approach to select which inputs from the unlabelled set should be labelled. Closer to our work, Bayesian PPI (Hofer et al., 2024) considers an alternative PPI estimator motivated by Bayesian ideas. However, their approach provides Bayesian credible intervals, which do not offer frequentist guarantees. Additionally, their approach achieves similar experimental performance to PPI, while we demonstrate that FAB-PPI may significantly improve upon PPI.
As discussed in Angelopoulos et al. (2023a, b), PPI has close ties with control variates for variance reduction (Glasserman, 2003, §4.1). In the case of mean estimation, the form of the PPI estimator is similar to the one proposed by Zhang et al. (2019). PPI is also related to work in semiparametric inference with missing data (Robins & Rotnitzky, 1995).
@@ 3.1 Control Variates
The method of control variates is a standard variance-reduction technique in Monte Carlo approximation (Glasserman, 2003, §4.1). For simplicity, we present the method in the scalar case, but extensions to the multivariate setting are available. Let $(Z,Y)$ be a pair of real-valued random variables, and assume we are interested in estimating $\mathbb{E}[Y]$ based on an iid sample $\{(Z_{i},Y_{i})\}_{i=1}^{n}$. Assuming $\mu=\mathbb{E}[Z]$ is known, one defines the control-variate estimator (CVE)
@@
where $\lambda\in\mathbb{R}$ is a tuning coefficient and $\overline{Z}$ and $\overline{Y}$ are the sample means of $(Z_{i})$ and $(Y_{i})$, respectively. The centred random variable $Z_{i}-\mu$ serves as a control variate to estimate $\mathbb{E}[Y]$. The CVE is a consistent and unbiased estimator of $\mathbb{E}[Y]$ with $\mathrm{var}(\widehat{Y}^{\mathrm{cv}}_{\lambda})=(\mathrm{var}(Y)-2\lambda\mathrm{cov}(Z,Y)+\lambda^{2}\mathrm{var}(Z))/n$, while $\mathrm{var}(\overline{Y})=\mathrm{var}(Y)/n$. Therefore, the CVE has smaller variance than $\overline{Y}$ whenever $\lambda\in(\min\{0,2\lambda^{\star}\},\,\max\{0,2\lambda^{\star}\})$, where the optimal coefficient is $\lambda^{\star}=\mathrm{cov}(Z,Y)/\mathrm{var}(Z)$. In this case, $\mathrm{var}(\widehat{Y}^{\mathrm{cv}}_{\lambda^{\star}})=(1-\rho_{Z,Y}^{2})\mathrm{var}(\overline{Y})$, where $\rho_{Z,Y}$ is the correlation between $Z$ and $Y$. The more correlated $Z$ and $Y$, the larger the variance reduction. By plugging the estimator
@@
as $n\to\infty$, where $s$ is the sample standard deviation of $\{(Y_{i}-\widehat{\lambda}Z_{i})\}_{i=1,\ldots,n}$. Hence, $\widehat{Y}^{\mathrm{cv}}_{\widehat{\lambda}}\pm z_{1-\alpha/2}s/\sqrt{n}$ is an asymptotically valid $(1-\alpha)$ CI for $\mathbb{E}[Y]$, whose asymptotic width is $2z_{1-\alpha/2}\sqrt{1-\rho_{Z,Y}^{2}}\sqrt{\mathrm{var}(Y)}/\sqrt{n}$.
@@ 3.2 Prediction-Powered Inference
PPI (Angelopoulos et al., 2023a) defines an estimator $\widehat{\theta}$ and a CI $\mathcal{C}_{\alpha}^{\mathrm{pp}}$ for a parameter of interest $\theta^{\star}$ satisfying Equation 2. In particular, let $\widehat{m}_{\theta}$ and $\widehat{\Delta}_{\theta}$ be some estimators of $m_{\theta}$ and $\Delta_{\theta}$. Using Equation 2, the estimator $\widehat{\theta}$ is defined as the solution, in $\theta$, to the equation
@@
where ++ denotes the Minkowski sum. Typical choices for $\widehat{m}_{\theta}$ and $\mathcal{T}_{\alpha-\delta}$ are the sample mean of the unlabelled data,
@@
Angelopoulos et al. (2023a) propose to use the sample mean
$\widehat{\Delta}^{\mathrm{\texttt{PP}}}_{\theta}$ $=\frac{1}{n}\sum_{i=1}^{n}\Big{(}\mathcal{L}_{\theta}^{\prime}(X_{i},Y_{i})-\mathcal{L}_{\theta}^{\prime}(X_{i},f(X_{i}))\Big{)}$ (10)
@@
Angelopoulos et al. (2023b) extend standard PPI by introducing an additional control-variate parameter $\lambda$, which they call power tuning parameter. The chosen $\widehat{m}_{\theta}$ is still the sample mean (9), while $\widehat{\Delta}^{\mathrm{\texttt{PP+}}}_{\theta}$ now takes the control variate form
@@
where $\widehat{\lambda}$ is estimated from the data. In this case, the centred control variate is $\mathcal{L}_{\theta}^{\prime}(X_{i},f(X_{i}))-\widehat{m}_{\theta}$, which depends only on the machine learning predictions. For the squared loss, we obtain
@@
While the definition of the PPI confidence interval (8) allows for merging any CIs $\mathcal{R}_{\delta}$ and $\mathcal{T}_{\alpha-\delta}$ for $\Delta_{\theta}$ and $m_{\theta}$, in practice these are often chosen to be CLT-based CIs that, once combined into $\mathcal{C}^{\mathrm{pp}}_{\alpha}$ give exact asymptotic coverage,
$\liminf_{n,N\to\infty}\mathrm{Pr}(\theta^{\star}\in\mathcal{C}^{\mathrm{pp}}_{\alpha})\geq 1-\alpha.$
@@ 3.3 Bayes-Optimal Confidence Regions
The FAB framework (Pratt, 1961, 1963; Yu & Hoff, 2018) aims to construct valid confidence regions with smaller expected volume. Let $W\mid\beta\sim\mathcal{N}(\beta,\sigma^{2})$ with some prior $\pi_{0}(\beta)$ and denote by $\pi(w)=\int p(w\mid\beta)\pi_{0}(\beta)d\beta$ the corresponding marginal likelihood. For $\alpha\in(0,1)$, let $\mathcal{C}_{\alpha}(w)$ be an exact $(1-\alpha)$ confidence region for $\beta$ based on the data $w$. That is, for any fixed $\beta_{0}$,
@@
Let $\mathrm{vol}(\mathcal{C}_{\alpha}(w))=\int_{\beta^{\prime}\in\mathcal{C}_{\alpha}(w)}d\beta^{\prime}$ be the volume of $\mathcal{C}_{\alpha}(w)$, and consider its expected value under the marginal likelihood $\pi$,
$\mathbb{E}[\mathrm{vol}(\mathcal{C}_{\alpha}(W))]=\int\mathrm{vol}(\mathcal{C}_{\alpha}(w))\pi(w)dw.$ (18)
@@ 4 FAB-PPI
As in PPI, we use the sample mean (9) as the estimator of $m_{\theta}$. For $\Delta_{\theta}$, we start by considering a consistent estimator $\widehat{\Delta}_{\theta}$, such as the sample mean (10) used in PPI, or the control variate estimator (12) used in PPI++. Throughout this section, we assume that Section 3.2 is satisfied. That is, a CLT holds for $\widehat{m}_{\theta}$ and $\widehat{\Delta}_{\theta}$ with respect to some estimators $(\widehat{\sigma}^{f}_{\theta})^{2}$ and $\widehat{\sigma}^{2}_{\theta}$ of $\mathrm{var}(\widehat{m}_{\theta})$ and $\mathrm{var}(\widehat{\Delta}_{\theta})$, respectively. In this setting, let $\pi_{0}(\Delta_{\theta};\tau_{n})$ be a prior on $\Delta_{\theta}$ with scale parameter $\tau_{n}$, which may depend on the labelled data through $\widehat{\sigma}_{\theta}$. Denote by $\ell(w;\sigma,\tau)$ the log-marginal likelihood, evaluated at $w$, of a Gaussian likelihood model with mean $\Delta$ and variance $\sigma^{2}$ under the prior $\pi_{0}(\Delta;\tau)$,
$\ell(w;\sigma,\tau)=\log\int_{\mathbb{R}}\mathcal{N}(w;\Delta,\sigma^{2})\pi_{0}(\Delta;\tau)d\Delta.$